OpenStudy (anonymous):

Eliminate the parameter. x = 3 cos t, y = 3 sin t

OpenStudy (anonymous):

So far I have: x/3 = cos(t) and y/3 = sin(t) then (y/3)^2 + (x/3)^2 = 1 but in the example in my lesson, it went a step further and squared out the demoninators which would make this answer: ((y^2)/9) + ((x^2)/9) = 1 is that what I should do for this? And then should I add them because of like terms? or no?

OpenStudy (anonymous):

The lesson said that it did that because it made it into the standard form of an ellipse. does this apply here, or is the regular answer correct?

OpenStudy (loser66):

correct

OpenStudy (anonymous):

So the answer would simply be: (y/3)^2 + (x/3)^2 = 1 right?

OpenStudy (loser66):

should be under the form of \[\dfrac{x^2}{9}+\dfrac{y^2}{9}=1\]

OpenStudy (anonymous):

oh, okay. so you do square out the denominators. Does this put this answer in the standard form of an ellipse?

OpenStudy (loser66):

if you define it is a ellipse, it is so. however, I would like to say it is a circle (x^2+y^2 =9)

OpenStudy (anonymous):

oh okay. My lesson just referred to it as an ellipse. Thanks!!